In producing chlorine through electrolysis, 100 watt power at 125 V is being consumed. How much chlorine per minute is liberated? E.C.E. of chlorine is 0.367 × 10-6 kg/coulomb.


        (a)    21.3 mg


        (b)    24.3 mg


        (c)    13.6 mg    


(d)    17.6 mg


Grade:upto college level

1 Answers

Aditi Chauhan
10 years ago

(d)

        Power = V × I

 

E.C.E. of chlorine is 0.367 × 10-6 kg/coulomb

 

Charge passing in one minute =48 coulomb

Chlorine precipitated = 0.367 × 10-6 × 48

     = 17.6 × 10-6 kg

     = 17.6mg